OpenStudy (iwanttogotostanford):

Simplify square root of negative 48.

OpenStudy (anonymous):

Can you think of any perfect squares that are factors of 48?

OpenStudy (iwanttogotostanford):

no not from the top of my head

OpenStudy (iwanttogotostanford):

its \[\sqrt{-48}\]

OpenStudy (anonymous):

I understand, but there's a method to my madness. You're going to have to simplify this radical, so you'll need to do this. The first few prefect squares are 1, 4, 9, 16, 25, 36, 49, etc. What is the largest of these that is a factor of 48?

OpenStudy (iwanttogotostanford):

4? because 4x12 is 48

OpenStudy (anonymous):

That's right, but is there a larger one?

OpenStudy (iwanttogotostanford):

I'm really rusty on my times table facts

OpenStudy (anonymous):

Use a calculator if it helps

OpenStudy (iwanttogotostanford):

ok

OpenStudy (iwanttogotostanford):

24

OpenStudy (anonymous):

Sorry, 24 is not a perfect square. I listed them above. What's the largest one that is a factor of 48?

OpenStudy (iwanttogotostanford):

I'm confused a bit, wouldn't it be 4? because 9x9 is 81 and thats too big

OpenStudy (anonymous):

9 is a perfect square because 3 x 3 = 9 16 is a perfect square because 4 x 4 = 16 25 is a perfect square because 5 x 5 = 25 etc. Which of those listed numbers is the largest one that is a factor of 48? You don't need to square them.

OpenStudy (iwanttogotostanford):

36

OpenStudy (anonymous):

Excellent. 16 x 3 = 48.

OpenStudy (iwanttogotostanford):

now what?

OpenStudy (anonymous):

Now, we're going to use the rules of working with radicals to simplify. Your question is\[\sqrt{-48}\]Having identified the largest perfect square that is a factor of 48 we can rewrite as follows\[\sqrt{-48}=\sqrt{\left( 16 \right)\left( -1 \right)\left( 3 \right)}\]Understand what we did here?

OpenStudy (iwanttogotostanford):

yes

OpenStudy (anonymous):

Good. Now the rules of radicals say that we can write this as follows\[\sqrt{-48} = \sqrt{\left( 16 \right)\left( -1 \right)\left( 3 \right)} = \sqrt{16}\sqrt{-1}\sqrt{3}\]You OK with that?

OpenStudy (iwanttogotostanford):

ok

OpenStudy (anonymous):

Good. You know what the square root of 16 is? And the square root of -1?

OpenStudy (iwanttogotostanford):

yes its 4 but i don't know the square root of -1

OpenStudy (anonymous):

You haven't studied imaginary numbers?

OpenStudy (iwanttogotostanford):

no, I'm learning them right now thats why i need help

OpenStudy (anonymous):

Well imaginary numbers are based on the square root of -1. It is an imaginary number that is given the symbol i. In other words\[\sqrt{-1} = i\]

OpenStudy (anonymous):

So, you have \(\sqrt{16} \sqrt{-1} \sqrt{3}\). And you know the square root of 16 and the square root of -1. Just substitute them in.

OpenStudy (iwanttogotostanford):

these are my answer choices negative 4 square root of 3 4 square root of negative 3 4 i square root of 3 4 square root of 3 i

OpenStudy (anonymous):

\[\sqrt{16}\sqrt{-1}\sqrt{3} = ?\]

OpenStudy (iwanttogotostanford):

would it be b or d?

OpenStudy (anonymous):

\(\sqrt{16}=4\) and \(\sqrt{-1} = i\) and \(\sqrt{3}\) can't be simplified any further. What does that give you?

OpenStudy (iwanttogotostanford):

D!?

OpenStudy (anonymous):

Well, the individual parts of D are correct but they're in the wrong order. Should have the rational number first, then the imaginary number i, then the radical. What other choice meets this description?

OpenStudy (iwanttogotostanford):

C then?

OpenStudy (anonymous):

That's correct. It's the convention that we write the answer \(4i\sqrt{3}\) rather than \(4\sqrt{3}i\) or \(\sqrt{3}i4\) or any other combination.

OpenStudy (iwanttogotostanford):

ok, thank you i was confused but now i understand better

OpenStudy (anonymous):

You're welcome